# Data Attribution for Text-to-Image Models by Unlearning Synthesized Images

### v1

@@ Abstract
The goal of data attribution for text-to-image models is to identify the training images that most influence the generation of a new image. We can define “influence” by saying that, for a given output, if a model is retrained from scratch without that output’s most influential images, the model should then fail to generate that output image. Unfortunately, directly searching for these influential images is computationally infeasible, since it would require repeatedly retraining from scratch. We propose a new approach that efficiently identifies highly-influential images. Specifically, we simulate unlearning the synthesized image, proposing a method to increase the training loss on the output image, without catastrophic forgetting of other, unrelated concepts. Then, we find training images that are forgotten by proxy, identifying ones with significant loss deviations after the unlearning process, and label these as influential. We evaluate our method with a computationally intensive but “gold-standard” retraining from scratch and demonstrate our method’s advantages over previous methods.
@@ 1 Introduction
Data attribution for text-to-image generation aims to identify which training images “influenced” a given output. The black-box nature of state-of-the-art image generation models [1, 2, 3, 4, 5, 6, 7, 8], together with the enormous datasets required [9], makes it extremely challenging to understand the contributions of individual training images. Although generative models can, at times, replicate training data [10, 11], they typically create samples distinct from any specific training image.
@@
We believe that a counterfactual definition of “influence” best matches the intuitive goal of attribution [12, 13]. Specifically, we say that a collection of training images is influential for a given output image if: removing those images from the training set and then retraining from scratch makes the model unable to generate the synthesized image. Unfortunately, directly searching for the most influential images according to this definition is computationally infeasible since it would require training an exponentially large number of new models from scratch.
@@
Hence, practical influence estimation requires effective approximations. For example, many approaches replace retraining with a closed-form approximation, computed separately for each training image [12, 14, 15, 13, 16]. For text-to-image attribution, these methods are outperformed by simple matching of off-the-shelf image features [17]. Wang et al. [18] use customization to study the effect of training a model towards an exemplar, but find limited generalization to the general large-scale training case. We aim for a tractable method that accurately predicts influence according to the counterfactual definition.
@@
We propose an approach to influence prediction with two key ideas (Figure 1). First, we can approximate removing a training image from a model by an optimization that we call unlearning. Unlearning increases the training loss of the target image while protecting unrelated concepts; we can then compute training loss for the original synthesized image. However, directly applying this idea would require unlearning separately for each training image. Our second main idea is to reverse the roles: we unlearn the synthesized image, and then evaluate which training images are represented worse by the new model. This requires only one unlearning optimization, rather than a separate unlearning for each training image.
@@
The methodology for unlearning is important. Unlearning a synthesized image by naively maximizing its loss leads to catastrophic forgetting [19], where the model fails to generate other unrelated concepts as well. Inspired by work on unlearning data for classifiers [20, 21], we mitigate this issue by regularizing gradient directions using Fisher information to retain pretrained information. Additionally, we find that updating only the key and value mappings in the cross-attention layers improves attribution performance. We show how “influence functions” [13, 15] can be understood as approximations to unlearning in Appendix A, but they are limited by their closed-form nature.
@@
  • We propose a novel method for data attribution for text-to-image models, unlearning the synthesized image and identifying which training images are forgotten.
@@
  • We rigorously show that our method is counterfactual predictive by omitting influential images, retraining, and checking that the synthesized image cannot be regenerated. Along with the existing Customized Model benchmark, we show our method identifies influential images more effectively than recent baselines based on customization and influence functions.
@@ 2 Related Work
Other methods explore different approaches. Inspired by the game-theory concept of Shapley value [29], several methods train models on subsets of training data and estimate the influence of a training point by comparing the models that had that training point in their data to those that did not [30, 31, 32, 33]. Pruthi et al. [34] estimate influence by tracking train-test image gradient similarity over the course of model training.
Recent methods perform attribution for diffusion-based image generation. Wang et al. [18] proposed attribution by model customization [35, 23], where a pretrained model is influenced by tuning towards an exemplar concept. Several works adapt TRAK [16], an influence function-based method, to diffusion models, extending it by attributing at specific denoising timesteps [12], or by improving gradient estimation and using Tikhonov regularization [14]. Unlike these methods, our method performs attribution by directly unlearning a synthesized image and tracking the effect on each training image. Our method outperforms existing methods in attributing both customized models and text-to-image models.
@@
Replication detection.   Shen et al. [50] identify repeated pictorial elements in art history. Somepalli et al. [11] and Carlini et al. [10] investigate text-to-image synthesis of perceptually-exact copies of training images. Unlike these works, our work focuses on data attribution for more general synthesis settings beyond replication.
@@ 3 Problem Setting and Evaluation
Our goal is to attribute a generated image to its training data. We represent the training data as $\mathcal{D}=\{({\bf x}_{i},{\bf c}_{i})\}_{i=1}^{N}$, where ${\bf x}\in\mathcal{X}$ denotes images and ${\bf c}$ represents the conditioning text. A learning algorithm $\mathcal{A}:\mathcal{D}\rightarrow\theta$ yields parameters of a generative model; for instance, $\theta=\mathcal{A}(\mathcal{D})$ is a model trained on $\mathcal{D}$. We focus on diffusion models that generate an image from a noise map $\epsilon\sim\mathcal{N}(0,{\bf I})$. A generated image from text c is represented as $\hat{\textbf{x}}=G_{\theta}(\epsilon,\textbf{c})$. To simplify notation, we write a text-image tuple as a single entity. A synthesized pair is denoted as $\hat{\mathbf{z}}=(\hat{\textbf{x}},\textbf{c})$, and a training pair is denoted as $\mathbf{z}_{i}=(\textbf{x}_{i},\textbf{c}_{i})\sim\mathcal{D}$. We denote the loss of an image x conditioned on c as $\mathcal{L}(\mathbf{z},\theta)$.
@@
Evaluating the model.   We then compare this “leave-$K$-out” model against $\theta_{0}=\mathcal{A}(\mathcal{D})$, the model trained with the entire dataset, and assess how much it loses its capability to represent $\hat{\mathbf{z}}$ in terms of both the loss change $\Delta\mathcal{L}(\hat{\mathbf{z}},\theta)$ and the capability to generate the same sample $\Delta G_{\theta}(\epsilon,\textbf{c})$.
@@
Second, if properly selected, the leave-$K$-out model should no longer be able to generate $\hat{\mathbf{x}}=G_{\theta}(\epsilon,\mathbf{c})$. For diffusion models, in particular, we can rely on the “seed consistency” property [12, 51, 52]. Georgiev et al. [12] find that images generated by two independently trained diffusion models from the same random noise $\epsilon$ have little variations. They leverage this property to evaluate attribution via $\Delta G_{\theta}(\epsilon,\textbf{c})$, the difference of generated images between $\theta_{0}$ and $\theta_{\hat{\mathbf{z}}}^{-K}$. An effective attribution algorithm should lead to a leave-$K$-out model generating images that deviate more from the original images, resulting in a larger $\Delta G_{\theta}(\epsilon,\textbf{c})$ value:
@@ 4 Attribution by Unlearning
In this section, we introduce our approach to attribute training images for a text-to-image model $\theta_{0}=\mathcal{A}(\mathcal{D})$, trained on dataset $\mathcal{D}$. Our goal is to find the highly influential images on a given synthetic image $\hat{\mathbf{z}}$ in dataset $\mathcal{D}$.
Formally, we define a data attribution algorithm $\boldsymbol{\tau}$, which given access to the training data, model, and learning algorithm, produces a set of scores, denoted by $\boldsymbol{\tau}(\hat{\mathbf{z}},\mathcal{D},\theta,\mathcal{A})\in\mathbb{R}^{N}$. The highest $K$ influencing images are then selected according to their scores. Due to the training set sizes $\mathcal{D}$ used in text-to-image models, we simplify the problem by individually estimating the influence of each training point: $\boldsymbol{\tau}(\hat{\mathbf{z}},\mathcal{D},\theta,\mathcal{A})=\left[\tau(\hat{\mathbf{z}},\mathbf{z}_{1}),\tau(\hat{\mathbf{z}},\mathbf{z}_{2}),\dots,\tau(\hat{\mathbf{z}},\mathbf{z}_{N})\right]$, where $\tau$ represents the factorized attribution function that estimates influence based on the synthesized sample and a training point. Although $\tau$ can access all training data, parameters, and the learning algorithm, we omit them for notational simplicity.
Given infinite compute, one can find critical sets of influential images by training a model from every possible subset of $K$ images from scratch and searching for one that “forgets” the synthesized image the most. However, the search space is combinatoric and infeasible in practice. To address this problem, we have made two modifications. First, rather than training from scratch, we use model unlearning—efficiently tuning a pretrained model to remove a data point. Second, rather than searching through the combinatoric space of training points, we instead apply unlearning to the synthesized image and then assess how effectively each training image is also forgotten as a result. As a heuristic to measure the degree of removal, we track the training loss changes for each training image after unlearning, and we find this effective for data attribution.
@@
where $\alpha$ controls the step size, $F^{-1}$ is the inverse of the Fisher information matrix. In practice, Newton updates allow us to achieve effective attribution with few iterations and, in some cases, as few as one step. We denote the unlearned model as $\theta_{-\hat{\mathbf{z}}}$. We provide details of the EWC loss and Newton update in Appendix A.
@@
Relation with influence functions.   Our method draws a parallel with the influence function, which aims to estimate the loss change of $\hat{\mathbf{z}}$ by removing a training point $\mathbf{z}$. However, training leave-one-out models on every training point is generally infeasible. Instead, the influence function relies on a heavier approximation to estimate the effect of perturbing a single training point, rather than actually forgetting the training samples. In contrast, our approach only requires running the unlearning algorithm once for a given synthesized image query. This allows us to use a more mild approximation and obtain a model that forgets the synthesized sample. Guo et al. [20] and Tanno et al. [21] explore a similar formulation for unlearning training images and draw a connection between their unlearning algorithms and influence function. Our approach aims to unlearn the synthesized image instead, which connects to influence function in a similar fashion. We discuss our method’s connection to influence function in more detail in Appendix A.3.
Optimizing a subset of weights.   To further regularize the unlearning process, we optimize a small subset of weights, specifically $W^{k}$ and $W^{v}$, the key and value projection matrices in cross-attention layers [56, 57]. In text-to-image models, cross-attention facilitates text-to-image binding, where $W^{k}$ identifies which features match each text token, while $W^{v}$ determines how to modify the features for the matched patches. We find that performing unlearning $W^{k}$ and $W^{v}$ is effective for attribution. Prior works also select the same set of parameters to improve fine-tuning [23] and unlearning [36].
@@ 5.1 Leave-$K$-out counterfactual evaluation
  • Influence function: Both TRAK and JourneyTRAK [12] are influence function-based methods that match the loss gradients of training and synthesized images, using random projection for efficiency. Both methods run the influence function on multiple models trained on the same dataset (20 in this test) and average the scores. The main difference is in the diffusion loss calculation: TRAK randomly samples and averages the loss over timesteps, while JourneyTRAK calculates it only at $t=400$ for synthesized images during counterfactual evaluation.
  • Random: We train models with $K$ random images removed, using 10 models per value of $K$, sweeping through $K=$ 500, 1000, 2000, 3000, 4000, and then 5000 to the full dataset size by increments of 2000.
For attribution methods, we use $K=500,1000,$ and $4000$, representing approximately $0.42\%$, $0.85\%$, and $3.4\%$ of the MSCOCO dataset, respectively. Densely sweeping $K$ is generally not feasible for attribution methods, as each queried synthesized image requires retraining a different set of $K$ images. To provide better intuition, we report the number of random images that need to be removed from the dataset to achieve the same forgetting effect as removing $K$ images from an attribution method.
@@
Tracking loss changes in leave-$K$-out models.   First, we report the change in DDPM loss for leave-$K$-out models in Figure 3. Matching in plain pixel or text feature space yields weak performance, while deep image features, particularly DINO, perform better. Interestingly, DINO outperforms influence function methods at $K=4000$, despite not being trained specifically for the attribution task. Fine-tuning image features with the Customized Model benchmark, such as CLIP (AbC), shows some improvement. However, in general, the improvement is limited, indicating that transferring from attributing customized models to general models remains challenging [18].
@@
Deviation of generated output in leave-$K$-out models.   Figure 4 shows the deviation in generated outputs for leave-$K$-out models, where all images are generated using the same noise input and text prompt. Our method yields the largest deviation with a small budget. DINO is the second strongest method in this evaluation, where it also starts significantly altering the generated output with higher $K$. In contrast, influence functions, including TRAK and JourneyTRAK, perform subpar in this test, whereas JourneyTRAK outperforms TRAK in this context. We report quantitative results, including MSE and CLIP similarities of images between the pre-trained and leave-$K$-out models, along with ablation studies and more qualitative results, in Appendix C.1.

### v2

@@ Abstract
The goal of data attribution for text-to-image models is to identify the training images that most influence the generation of a new image. Influence is defined such that, for a given output, if a model is retrained from scratch without the most influential images, the model would fail to reproduce the same output. Unfortunately, directly searching for these influential images is computationally infeasible, since it would require repeatedly retraining models from scratch. In our work, we propose an efficient data attribution method by simulating unlearning the synthesized image. We achieve this by increasing the training loss on the output image, without catastrophic forgetting of other, unrelated concepts. We then identify training images with significant loss deviations after the unlearning process and label these as influential. We evaluate our method with a computationally intensive but “gold-standard” retraining from scratch and demonstrate our method’s advantages over previous methods.
@@ 1 Introduction
Data attribution for text-to-image generation aims to identify which training images “influenced” a given output. The black-box nature of modern image generation models [1, 2, 3, 4, 5, 6, 7, 8], together with the enormous datasets required [9], makes it extremely challenging to understand the contributions of individual training images. Although generative models can, at times, replicate training data [10, 11], they typically create samples distinct from any specific training image.
@@
We believe that a counterfactual definition of “influence” best matches the intuitive goal of attribution [12, 13]. Specifically, a collection of training images is influential for a given output image if removing those images from the training set and then retraining from scratch makes the model unable to reproduce the same synthesized image. Unfortunately, directly searching for the most influential images according to this definition is computationally infeasible since it would require training an exponentially large number of new models from scratch.
@@
Hence, practical influence estimation requires effective approximations. For example, many approaches replace retraining with a closed-form approximation, computed separately for each training image [12, 14, 15, 13, 16]. For text-to-image attribution, these methods are outperformed by simple matching of off-the-shelf image features [17]. Wang et al. [18] use model customization [19] to study the effect of training a model towards an exemplar, but find limited generalization to the general large-scale training case. We aim for a tractable method that accurately predicts influence according to the counterfactual definition.
@@
We propose an influence prediction approach with two key ideas (Figure 1). First, we approximate the removal of a training image from a model through an optimization procedure termed unlearning [20, 21, 22], which increases the training loss of the target image while preserving unrelated concepts. We then compute the training loss for the original synthesized image. However, directly applying this idea would require unlearning separately for every single training image, which is also costly. Our second main idea is to reverse the roles: we unlearn the synthesized image, and then evaluate which training images are represented worse by the new model. This method requires only one unlearning optimization, rather than a separate unlearning step for each training image.
@@
The methodology for unlearning is important. Unlearning a synthesized image by naively maximizing its loss leads to catastrophic forgetting [23], where the model also fails to generate other unrelated concepts. Inspired by work on unlearning data for classifiers [20, 24], we mitigate this issue by regularizing gradient directions using the Fisher information to retain pretrained information. Additionally, we find that updating only the key and value mappings in the cross-attention layers improves attribution performance. We show how “influence functions” [13, 15] can be understood as approximations to unlearning in Appendix A, but they are also limited by their closed-form approximation nature.
@@
  • We propose a new data attribution method for text-to-image models, by unlearning the synthesized image and identifying which training images are forgotten.
@@
  • We rigorously show that our method is counterfactual predictive by omitting influential images, retraining, and checking that the synthesized image cannot be regenerated. Along with the existing Customized Model benchmark, our method identifies influential images more effectively than recent methods based on customization and influence functions.
@@ 2 Related Work
Other methods explore different approaches. Inspired by the game-theory concept of Shapley value [31], several methods train models on subsets of training data and estimate the influence of a training point by comparing the models with and without that training point  [32, 33, 34, 35]. Pruthi et al. [36] estimate influence by tracking train-test image gradient similarity throughout model training.
Recent methods have started tackling attribution for diffusion-based image generation. Wang et al. [18] proposed attribution by model customization [37, 19], where a pretrained model is influenced by tuning towards an exemplar concept. Several works adapt TRAK [16], an influence function-based method, to diffusion models, extending it by attributing at specific denoising timesteps [12], or by improving gradient estimation and using Tikhonov regularization [14]. Unlike these methods, our method performs attribution by directly unlearning a synthesized image and tracking the effect on each training image. Our method outperforms existing methods in attributing both customized models and text-to-image models. Concurrent works also apply machine unlearning for attribution tasks [38, 39]. Unlike their approaches, we find that applying strong regularization during unlearning is crucial to obtaining good performance in our tasks.
@@
Replication detection.   Shen et al. [52] identify repeated pictorial elements in art history. Somepalli et al. [11] and Carlini et al. [10] investigate the text-to-image synthesis of perceptually-exact copies of training images. Unlike these, our work focuses on data attribution for more general synthesis settings beyond replication.
@@ 3 Problem Setting and Evaluation
Our goal is to attribute a generated image to its training data. We represent the training data as $\mathcal{D}=\{({\bf x}_{i},{\bf c}_{i})\}_{i=1}^{N}$, where ${\bf x}\in\mathcal{X}$ denotes an image and ${\bf c}$ represents its conditioning text. A learning algorithm $\mathcal{A}:\mathcal{D}\rightarrow\theta$ yields parameters of a generative model; for instance, $\theta=\mathcal{A}(\mathcal{D})$ is a model trained on $\mathcal{D}$. We focus on diffusion models that generate an image from a noise map $\epsilon\sim\mathcal{N}(0,{\bf I})$. A generated image from text c is represented as $\hat{\textbf{x}}=G_{\theta}(\epsilon,\textbf{c})$. To simplify notation, we write a text-image tuple as a single entity. A synthesized pair is denoted as $\hat{\mathbf{z}}=(\hat{\textbf{x}},\textbf{c})$, and a training pair is denoted as $\mathbf{z}_{i}=(\textbf{x}_{i},\textbf{c}_{i})\sim\mathcal{D}$. We denote the loss of an image x conditioned on c as $\mathcal{L}(\mathbf{z},\theta)$.
@@
Evaluating the model.   We then compare this “leave-$K$-out” model against $\theta_{0}=\mathcal{A}(\mathcal{D})$, the model trained with the entire dataset, and assess how much it loses its capability to represent $\hat{\mathbf{z}}$ in terms of both the loss change $\Delta\mathcal{L}(\hat{\mathbf{z}},\theta)$ and the capability to generate the same sample $\Delta G_{\theta}(\epsilon,\textbf{c})$ from the same input noise $\epsilon$ and text c.
@@
Second, if properly selected, the leave-$K$-out model should no longer be able to generate $\hat{\mathbf{x}}=G_{\theta}(\epsilon,\mathbf{c})$. For diffusion models, we can particularly rely on the “seed consistency” property [12, 53, 54]. Georgiev et al. [12] find that images generated from the same random noise have little variations, even when generated by two independently trained diffusion models on the same dataset. They leverage this property to evaluate attribution via $\Delta G_{\theta}(\epsilon,\textbf{c})$, the difference of generated images between $\theta_{0}$ and $\theta_{\hat{\mathbf{z}}}^{-K}$. An effective attribution algorithm should lead to a leave-$K$-out model generating images that deviate more from the original images, resulting in a larger $\Delta G_{\theta}(\epsilon,\textbf{c})$ value:
@@
Choice of loss $\mathcal{L}(\mathbf{z},\theta)$.   We focus on DDPM loss introduced by Ho et al. [56], the standard loss used to train diffusion models. Diffusion models learn to predict the noise added to a noisy image $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, where $\mathbf{x}_{0}$ is the original image, $\epsilon$ is the Gaussian noise, $\bar{\alpha}_{t}$ and $t$ controls the noise strength. $t$ is an integer timestep sampled between 1 to $T$, where $T$ is typically set to 1000. A larger $t$ implies more noise added. DDPM loss optimizes for the noise prediction task: $\mathbb{E}[\|\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{c},t)-\epsilon\|^{2}],$ where $\epsilon_{\theta}(\cdot)$ is the denoiser for noise prediction, and $c$ denotes text condition.
@@ 4 Attribution by Unlearning
In this section, we introduce our attribution approach for a text-to-image model $\theta_{0}=\mathcal{A}(\mathcal{D})$, trained on dataset $\mathcal{D}$. We aim to find the highly influential images on a given synthetic image $\hat{\mathbf{z}}$ in dataset $\mathcal{D}$.
If we had infinite computes and a fixed set of $K$ images, we could search for every possible subset of $K$ images and train models from scratch. The subset whose removal leads to the most ’‘forgetting” of the synthesized image would be considered the most influential. Of course, such a combinatoric search is impractical, so we simplify the problem by estimating the influence score of each training point individually and selecting the top $K$ influential images based on their scores.
Formally, we define a data attribution algorithm $\boldsymbol{\tau}$, which given access to the training data, model, and learning algorithm, estimates the influence of each training point, denoted by $\boldsymbol{\tau}(\hat{\mathbf{z}},\mathcal{D},\theta,\mathcal{A})=\left[\tau(\hat{\mathbf{z}},\mathbf{z}_{1}),\tau(\hat{\mathbf{z}},\mathbf{z}_{2}),\dots,\tau(\hat{\mathbf{z}},\mathbf{z}_{N})\right]$. $\tau$ represents the factorized attribution function that estimates influence based on the synthesized sample and a training point. Although $\tau$ can access all training data, parameters, and the learning algorithm, we omit them for notational simplicity.
One potential algorithm for $\tau(\hat{\mathbf{z}},\mathbf{z})$ is to remove $\mathbf{z}$ from the training set, retrain a model and check its performance on $\hat{\mathbf{z}}$. However, this method is infeasible due to the size of the training set. To overcome this, we have introduced two key ideas. First, rather than training from scratch, we use model unlearning—efficiently tuning a pretrained model to remove a data point. Second, rather than unlearning training points, we apply unlearning to the synthesized image and assess how effectively each training image is forgotten. To measure the degree of removal, we track the training loss changes for each training image after unlearning, and we find this effective for data attribution.
@@
where $\alpha$ controls the step size, $F^{-1}$ is the inverse of the Fisher information matrix. In practice, Newton updates allow us to achieve effective attribution with few iterations and, in some cases, as few as one step. We denote the unlearned model as $\theta_{-\hat{\mathbf{z}}}$. We provide details of the EWC loss, and Newton update in Appendix A.
@@
Relation with influence functions.   Our method draws a parallel with the influence function, which aims to estimate the loss change of $\hat{\mathbf{z}}$ by removing a training point $\mathbf{z}$. However, training leave-one-out models for every training point is generally infeasible. Instead, the influence function relies on a heavier approximation to estimate the effect of perturbing a single training point, rather than actually forgetting the training samples. In contrast, our approach only requires running the unlearning algorithm once for a synthesized image query. This allows us to use a milder approximation and obtain a model that forgets the synthesized sample. Guo et al. [20] and Tanno et al. [24] explore a similar formulation for unlearning training images and draw a connection between their unlearning algorithms and influence function. Our approach aims to unlearn the synthesized image instead, which connects to influence function in a similar fashion. We discuss our method’s connection to influence function in more detail in Appendix A.3.
Optimizing a subset of weights.   To further regularize the unlearning process, we optimize a small subset of weights, specifically $W^{k}$ and $W^{v}$, the key and value projection matrices in cross-attention layers [59, 60]. In text-to-image models, cross-attention facilitates text-to-image binding, where $W^{k}$ identifies which features match each text token, while $W^{v}$ determines how to modify the features for the matched patches. We observe that performing unlearning $W^{k}$ and $W^{v}$ is effective for attribution. Prior works also select the same set of parameters to improve fine-tuning [19] and unlearning [40].
@@ 5.1 Leave-$K$-out counterfactual evaluation
  • Random: We train models with $K$ random images removed, using 10 models per value of $K$.
@@
  • DataInf [62]: an influence estimation method based on approximating matrix inverses.
  • TRAK [16] and JourneyTRAK [12] are influence function-based methods that match the loss gradients of training and synthesized images, using random projection for efficiency. Both methods run the influence function on multiple models trained on the same dataset (20 in this test) and average the scores. The main difference is in the diffusion loss calculation: TRAK randomly samples and averages the loss over timesteps, while JourneyTRAK calculates it only at $t=400$ for synthesized images during counterfactual evaluation.
  • D-TRAK [14]: a concurrent work that extends TRAK by changing the denoising loss function into a square loss during influence computation. As mentioned by the authors, D-TRAK yields unexpectedly stronger performance even though the design choice is not theoretically understood. Different from TRAK, D-TRAK yields competitive performance with a single model.
For attribution methods, we use $K=500,1000,$ and $4000$, representing approximately $0.42\%$, $0.85\%$, and $3.4\%$ of the MSCOCO dataset, respectively.
@@
K=500 K=1000 K=4000 K=500 K=1000 K=4000 K=500 K=1000 K=4000
Random Random 3.5±0.03 3.5±0.03 3.5±0.03 4.1±0.06 4.1±0.06 4.0±0.06 7.9±0.03 7.8±0.03 7.9±0.03
@@
[…]
CLIP 4.2±0.14 4.7±0.18 6.4±0.32 4.4±0.20 4.6±0.21 5.2±0.22 7.6±0.09 7.5±0.08 6.8±0.08
DINO 4.8±0.15 5.6±0.20 8.1±0.35 4.5±0.16 5.3±0.22 5.9±0.21 7.4±0.09 7.1±0.09 6.3±0.10
[…]
DINO (AbC) 4.8±0.15 5.5±0.20 8.1±0.35 4.8±0.22 4.9±0.20 5.8±0.21 7.5±0.09 7.2±0.09 6.3±0.09
[…]
TRAK 5.2±0.14 5.8±0.16 7.1±0.16 4.7±0.21 4.7±0.24 4.7±0.20 7.6±0.09 7.6±0.09 7.5±0.09
JourneyTRAK 4.4±0.11 4.9±0.13 5.7±0.15 4.8±0.19 5.4±0.23 5.4±0.24 7.7±0.09 7.5±0.09 7.5±0.09
D-TRAK 5.4±0.16 6.6±0.22 9.6±0.33 5.9±0.24 6.4±0.25 7.8±0.30 7.3±0.10 7.1±0.09 6.4±0.09
[…]
Table 1: Leave-$K$-out baseline comparisons. Given a synthesized image $\hat{\mathbf{z}}$, we train leave-$K$-out models for each of the attribution methods and track $\Delta\mathcal{L}(\hat{\mathbf{z}},\theta)$, the increase in loss change, and $\Delta G_{\theta}(\epsilon,\mathbf{c})$, deviation of generation. We report results over 110 samples, and gray shows the standard error. Bolded and underlined are the best and second best performing method, respectively.
Tracking loss changes in leave-$K$-out models.   First, we report the change in DDPM loss for leave-$K$-out models in Table 1. Matching in plain pixel or text feature space yields weak performance, while deep image features, particularly DINO, perform better. Interestingly, DINO outperforms most influence function methods at $K=4000$, despite not being trained specifically for the attribution task. Fine-tuning image features with the Customized Model benchmark, such as CLIP (AbC), shows some improvement. However, in general, the improvement is limited, indicating that transferring from attributing customized models to general models remains challenging [18].
Among influence functions, DataInf performs poorly. According to the paper, the matrix inverse approximation DataInf uses is more suitable for LoRA fine-tuned models rather than text-to-image models trained from scratch. This approximation leads to poor performance.
TRAK significantly outperforms JourneyTRAK. We hypothesize that this is because JourneyTRAK collects gradients only for denoising loss at timestep $t=400$, making it less effective for identifying influential images that affect the DDPM training loss across different noise levels. On the other hand, D-TRAK is the best performing influence-function-based method, although the reason behind its strong performance is not well-understood [14].
Our method consistently performs best across all $K$ values, outperforming both influence functions and feature-matching methods.
[…]
Deviation of generated output in leave-$K$-out models.   Figure 3 and Table 1 shows the deviation in generated outputs for leave-$K$-out models, where all images are generated using the same noise input and text prompt. Consistent with the loss change evaluation, D-TRAK, DINO, and our method yield the largest deviation with a small budget. While the three methods perform similarly well in CLIP similarity, D-TRAK outperforms our method in MSE deviations. In contrast, TRAK and JourneyTRAK have formulations similar to D-TRAK, but they perform subpar in this test. Interestingly, while D-TRAK yields significant performance gain by changing the loss function, we did not observe the same improvement when applying the same loss function changes to our method. We discuss this more in Appendix C.1..
In addition, we include analysis on whether unlearned models and leave-$K$-out models forget only the specific concept, along with more qualitative results, in Appendix C.1.
Ablation studies.   We study two design choices: (1) the effect of EWC regularization and (2) different weight subsets for optimization. Table 2 shows the results. We find that unlearning without EWC loss greatly hurts attribution performance, indicating the importance of regulating unlearning with Fisher information. We also find that using a subset of weights to unlearn leads to better attribution in general. We test three weight subset selection schemes (Attn, Cross Attn, Cross Attn KV), all of which outperform the version using all weights. Among them, updating Cross Attn KV performs the best, consistent with findings from model customization [19, 63] and unlearning [40].
[…]
Variation EWC? K=500 K=1000 K=4000 K=500 K=1000 K=4000 K=500 K=1000 K=4000
SGD (1 step) 3.5±0.10 3.5±0.10 3.5±0.11 3.9±0.16 3.9±0.17 4.1±0.20 7.9±0.09 7.8±0.09 7.8±0.09
SGD (10 step) 3.4±0.10 3.5±0.10 3.5±0.10 3.7±0.16 3.9±0.17 3.9±0.15 7.9±0.09 7.9±0.09 7.8±0.09
[…]
Attn 5.5±0.17 6.6±0.23 9.3±0.33 5.1±0.23 5.3±0.22 6.2±0.22 7.4±0.10 7.2±0.09 6.7±0.09
Cross Attn 5.5±0.17 6.5±0.24 9.1±0.33 4.8±0.20 5.3±0.22 6.0±0.23 7.3±0.10 7.2±0.10 6.6±0.09
Cross Attn KV 5.6±0.16 6.7±0.22 9.8±0.32 5.1±0.21 5.7±0.24 6.1±0.22 7.3±0.09 7.0±0.09 6.4±0.10
Table 2: Leave-$K$-out ablation studies. We ablate our design choices and report $\Delta\mathcal{L}(\hat{\mathbf{z}},\theta)$ and $\Delta G_{\theta}(\epsilon,\mathbf{c})$ as in Table 1. We find that naive unlearning (SGD without EWC regularization) is not effective. We then compare four different sets of weights to unlearn and find that cross-attention $W^{k}$, $W^{v}$ (Cross Attn KV) outperforms other configurations. Bolded are the best performing method, and gray shows the standard error.
@@
5.2 Customized Model Benchmark
Wang et al. [18] focus on a specialized form of attribution: attributing customized models trained on an individual or a few exemplar images. This approach provides ground truth attribution, since the images generated by customized models are computationally influenced by exemplar images. While this evaluation has limited generalization to attribution performance with larger training sets, it is the only tractable evaluation for attributing large-scale text-to-image models to date.
Evaluation protocol.   Since the Customized Model Benchmark has ground truth, the problem is evaluated as a retrieval task. We report Recall@K and mAP, measuring the success of retrieving the exemplar images amongst a set including 100K LAION images. We compare with Wang et al.’s feature-matching approach that finetunes on the Customized Model dataset, referred to as DINO (AbC) and CLIP (AbC). We also compare with D-TRAK, the best-performing influence function method in our previous MSCOCO experiment. For our evaluation, we selected a subset of the dataset comprising 20 models: 10 object-centric and 10 artist-style models. We select 20 synthesized images with different prompts for each model, resulting in 400 synthesized image queries.
Comparing with other methods.   We report Recall@10 and mAP in Figure 6. Our method performs on par with baselines for object-centric models, while significantly outperforming them on artist-style models. Although CLIP (AbC) and DINO (AbC) are fine-tuned for this attribution task, the feature matching approach can sometimes confuse whether to attribute a synthesized image to style-related or object-related images. In contrast, our method, which has access to the model itself, traces influential training images more effectively. D-TRAK performs worse than our method despite also having model access, which suggests that the influence approximations could be less accurate for larger models. In Figure 5, we show a qualitative example. While DINO (AbC) and CLIP (AbC) can retrieve visually or semantically similar images, our method successfully identifies the exemplars in both cases. We include ablation studies in Appendix C.2.
Figure 5: Qualitative examples on the Customized Model benchmark. The red boxes indicate ground truth exemplar images used for customizing the model. Both our method and AbC baselines successfully identify the exemplar images on object-centric models (left), while our method outperforms the baselines with artistic style models (right).